Translate to a system of equations and solve. (Enter your answers as a comma-separated list.)
The sum of two numbers is 15. One number is 3 less than the other. Find the numbers.

Answer by Boreal(15235)

x+y=15
x=y-3
substitute
y-3+y=15
2y=18
y=9
y-3=6
the numbers are 6 and 9.